The elementary reaction

2
H
2
O
(
g
)




2
H
2
(
g
)
+
O
2
(
g
)
2
H
2
O
(
g
)




2
H
2
(
g
)
+
O
2
(
g
)

proceeds at a certain temperature until the partial pressures of
H
2
O
H
2
O
,
H
2
H
2
, and
O
2
O
2
reach
0.063
bar
0.063
bar
,
0.0024
bar
0.0024
bar
, and
0.0017
bar
0.0017
bar
, respectively, at equilibrium. What is the value of the equilibrium constant at this temperature
1
Expert's answer
2020-04-06T11:46:19-0400

"K_p=\frac{p_{H_2}^2\times p_{O_2}}{p_{H_2O}^2}=\frac{0.0024^2\times0.0017}{0.063^2}=0.2467\times10^{-5}bar"